3 ∫ 2 (2x 2 + 1) dx

32 (2x2 + 1) dx

Grade:12

2 Answers

Anusha
9 years ago
Ans:- 32 (2x2 +1)dx
=2[(x2+1 /2+1)]23+[x]23
=2[(x3 /3)]23 +[x]23
=2[(33 /3)+(23 /3)]+[3-2]
=2[(27+8)/3]+1
=2[(35/3)]+1
=(70/3)+1
=73/3
Miryala Gopalakrishna
9 years ago
integration of (2(X^2)+1)dx from 2 to 3 is equal to
(2((x^3)/3)+x)  from 2 to3
=(2((3^3)/3)+3)-(2((2^3)/3)+2)
=(2*(27/3)+3)-(2*(8/3)+2)
=(18+3)-((16/3)+2)
=(21)-(22/3)
=41/3